HIGH RESOLUTION ROVIBRATIONAL SPECTRA OF FIVE-MEMBERED RINGS WITH HET-EROATOMS: NEW RESULTS FROM THE FAR-INFRARED BEAMLINE OF THE CANADIAN LIGHT SOURCE

D. W. TOKARYK, Department of Physics and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, Canada E3B 5A3; \underline{J. A. VAN WIJNGAARDEN}, Department of Chemistry, University of Manitoba, Winnipeg, MB, Canada R3T 2N2.

The ground state spectra and structures of five-membered ring molecules like pyrrole, furan and thiophene are well established via microwave spectroscopy. Some rotationally-resolved studies of the vibrational bands of these molecules have previously been conducted at high resolution by Fourier transform and infrared diode laser spectroscopy; however, the data obtained via Fourier transform spectroscopy suffers in quality due to the low light flux of conventional light sources (particularly at frequencies lower than 600 cm^{−1}), and due to the requirement that the entrance iris be very small to obtain spectra at high resolution. We report new rotationally-resolved vibrational spectra of five-membered ring molecules using the intense, spatially-confined infrared light produced by the Canadian Light Source synchrotron, which we have coupled into a Bruker IFS125 Fourier transform spectrometer. Spectra of exceptionally high quality were taken at the full resolution of the instrument (0.001 cm^{−1}), both with the synchrotron source (400-700 cm^{−1}) and the internal globar source (700-1000 cm^{−1}). Comparisons of the 400-700 cm^{−1} spectra with those taken with a conventional globar source are given, and analysis of the rotational structure of these bands is presented.